# Deep Reinforcement Learning for Traveling Purchaser Problems

### v5

@@ 3.2.3 Transition
where $q_{v^{t}k}$ is the available quantity of product $k$ supplied at market $v^{t}$. Otherwise, if the next node $v^{t}$ is the depot, indicating a complete route is formed, the MDP terminates and then the reward calculation procedure for the terminal state is executed. Note that the TPP instance $U$, that is, the static part of the states, remains fixed throughout the state transitions in an MDP.
@@ 3.2.4 Reward
After the sequential route construction process described above, the second stage, i.e., purchase planning, is executed, which is integrated into the reward calculation procedure for the terminal state. Given the constructed route, a linear transportation problem is solved to derive the optimal purchasing plan for the visited markets. In the transportation problem, each source point corresponds to a market on the route, and each destination point corresponds to a product to be purchased. The cost to purchase a unit of product $k$ from market $i$ is exactly the unit transportation cost from source point $i$ to destination point $k$ in the transportation problem. This transportation problem can be readily solved using an off-the-shelf LP solver or a polynomial-time dynamic programming algorithm (Munkres, 1957). So far, a complete solution, including a route and the associated purchasing plan, has been obtained.
The objective of TPPs is to minimize the sum of traveling and purchasing costs while meeting all problem constraints. Since the feasibility of the route and the purchasing plan have been guaranteed by the masking rules (Section 3.2.2), we define the reward for the terminal state as the negative objective value, i.e., the sum of traveling and purchasing costs, and set zero-reward for all intermediate states. The reward is used to update the policy network through a DRL training algorithm, which will be specifically described later in Section 5.2. This way, the policy network is trained to optimize the global objective, that is, to construct a high-quality route that minimizes not only the traveling cost itself but also the subsequent purchasing cost based on it, thus to minimize the total cost.
@@ 3.2.5 Policy
Given the state $s_{t}$ at each step $t$, the action $a_{t}$ is selected according to a stochastic policy $p_{\theta}(a_{t}\mid s_{t})$, which can be seen as an action distribution over the action set $A_{t}$. Since the transition from a state $s_{t}$ to the next state $s_{t+1}$ is deterministic given an action $a_{t}$ (Section 3.2.3), the joint probability to sample a route $\pi$ based on the stochastic policy $p_{\theta}(a_{t}\mid s_{t})$ can be factorized according to the chain rule as:
@@
where the route is constructed in $T$ steps, and the action $a_{t}$ is sampled based on the policy $p_{\theta}(a_{t}\mid s_{t})$ at each step. In our DRL-based approach, the policy is parameterized as a deep neural network $\theta$, where the input of the policy network is the state $s_{t}$ and the output is the policy, i.e., the action distribution $p_{\theta}(a_{t}\mid s_{t})$. It can be seen that in our framework, the policy network plays a determinative role in the quality of solutions. Our training goal is to learn an effective policy network that can produce high-reward solutions with high probabilities and low-reward solutions with low probabilities.
@@ 4 Bipartite Graph Representation for TPPs
Similar to existing DRL-based methods for routing problems (Kool et al., 2019; Fu et al., 2021), a natural and straightforward idea is to represent a TPP instance as a complete or k-NN graph, where each market is represented as a node, which is connected to each other, and the product supply information at each market are represented as node features. However, despite containing necessary information for the TPP instance, such representations impose many drawbacks when used in DRL. First, the feature dimensions vary with the number of products in the TPP instance being solved. For example, considering two features for each product: the available quantity and price, a 50-product instance requires a 100-dimensional feature vector for each node to record these 50 products, while an instance with 100 products would require 200 dimensions. This will lead to a computational mismatch in the forward-propagation of neural networks. In other words, a policy network can only be applied on fixed-sized TPP instances, unable to adapt to instances with different sizes even if only one additional product is included, which significantly limits its practical usability. Second, the relations between markets and products, such as the substitutions and complementarities in product supply between different markets, are very important for constructing a low-cost route, but such information is largely ignored in the aforementioned complete or k-NN graph representations. To address these limitations, we design a novel bipartite graph representation, which is built upon the MILP formulation of TPPs (Section 2).
@@
Bipartite graphs have been demonstrated effective in representing general MILP problems (Nair et al., 2020; Morabit et al., 2021), where the variables and constraints are represented as two sets of nodes, and an edge connects a variable node and a constraint node if the corresponding column contributes to this constraint. Readers are referred to Chen et al. (2023) for a more detailed introduction and analysis. However, directly applying this representation to the MILP formulation (1)-(9) for TPPs can introduce unnecessary redundancy: the parameters to characterize a specific TPP instance exist only in the objective function (1) and constraints (2)-(3), whereas constraints (4)-(6), which are used to form a simple cycle, have already been guaranteed by the masking rules (Section 3.2.2) and therefore need not be explicitly represented for the policy network. Moreover, the constraints and variables associated with the same element (e.g., a market or a product) can be further merged to reduce the graph size and thus improve the computational efficiency of the policy network.
@@
Therefore, we design a bipartite graph representation for TPPs based on the objective function (1) and constraints (2)-(3), with appropriate mergence for the constraints and variables therein. Our bipartite graph representation is shown as Figure 2. We represent the constraints associated only with products (i.e., constraints (2)) and the variables associated only with markets/depot (i.e., variables $v_{i}$ and $x_{ij}$) as two sets of nodes, and the constraints and variables associated with both products and markets (i.e., constraints (3) and variables $z_{ik}$) as edges between the product nodes and market/depot nodes. Specifically, each variable $v_{i}$ is represented as a market/depot node, and the objective coefficients $c_{ij}$ of variables $x_{ij}$ are represented as coordinate features of the market/depot nodes. Each constraint in (2) is represented as a product node, with the right-hand side values $d_{k}$ attached as node features. An edge connects a market node $i$ and a product node $k$ if the associated variable $z_{ik}$ is not always restricted to 00 in constraints (3), that is, product $k$ is available at market $i$. The edge feature is a 2-dimensional vector, which consists of the corresponding objective coefficient $p_{ik}$ of variable $z_{ik}$ and the constraint coefficient $q_{ik}$ in constraints (3). For simplicity in the remainder of this paper, we use the term “market nodes” to refer to “market/depot nodes” if it does not cause confusion. We remark that our bipartite graph representation can be viewed from a more intuitive perspective: the two core elements in TPPs, i.e., markets and products, are represented as two sets of nodes, and their connections, i.e., available quantities and prices of the products at the markets, are characterized using the edges between them.
@@
We emphasize that our bipartite graph representation effectively addresses the limitations of the complete or k-NN graph representations mentioned above. First, the node and edge feature dimensions are invariant to the number of markets and products, allowing the design of a size-agnostic policy network, such that it can be flexibly adapted to TPP instances of varying sizes. Second, the bipartite graph explicitly characterizes the relations between markets and products through the edges connecting market nodes and product nodes. The message-passing along edges can effectively capture the relational information between markets and products, which is useful for the decision tasks.
@@ 5 Policy Network
In this section, we introduce the architecture of our policy network, which is used in the first stage for route construction. In addition, at the end of this section, we will describe the basic training algorithm for updating the policy network during training.
@@ 5.1 Policy Network Architecture
As aforementioned, at each decision step $t$, the policy network $\theta$ takes the state $s_{t}$ (consisting of the TPP instance $U$ being solved, which is represented as a bipartite graph, the partial route $\pi^{t}$, and the remaining demand $d^{t}$) as input, and outputs a distribution $p_{\theta}(a_{t}|s_{t})$ over the actions $a_{t}\in A_{t}$ to decide the next node to visit. The architecture of our policy network is illustrated as Figure 3. The policy network is composed of 1) an input embedding module, 2) a market encoder, and 3) a decoder.
The encoding process is designed based on the structure of bipartite graph. Specifically, the input embedding module takes the bipartite graph representation of instance $U$ as input, performing message-passing along edges and producing high-dimensional embeddings for the market and product nodes. Then, the market node embeddings are processed through the market encoder to extract further information for the TPP instance. After the encoding process, the decoder is executed iteratively to construct the route in a sequential manner. Specifically, at each decision step $t$, the decoder receives a decoding context, which contains the embeddings of the bipartite graph and the dynamic part of current state $s_{t}$ (i.e., $\pi^{t}$ and $d^{t}$) as input. It then outputs a distribution $p_{\theta}(a_{t}\mid s_{t})$, i.e., the policy, from which an action $a_{t}$ is selected, determining the next node to visit, and the dynamic part $\pi^{t}$ and $d^{t}$ of the state are updated to start the next decoding step. Note that given a TPP instance $U$, the bipartite graph representation of $U$ is fixed throughout the state transitions in a complete MDP. Therefore, we execute the input embedding module and market encoder only once in the initial state, and then execute the decoder iteratively with the decision steps of the MDP. In the following, we will present a detailed introduction to the three components of our policy network.
Figure 3: The architecture of our policy network. The policy network is composed of 1) an input embedding module, 2) a market encoder, and 3) a decoder. The bipartite graph is embedded through the input embedding module and the market encoder. Then, the decoder is executed iteratively: at each step $t$, the decoder takes as input a decoding context, and outputs an action distribution $p_{\theta}(a_{t}\mid s_{t})$, from which an action $a_{t}$ is selected, determining the next node to visit.
@@ 5.1.1 Input Embedding Module
Taking the bipartite graph representation of instance $U$ as input, the input embedding module exploits the raw features and topology of the bipartite graph to produce a high-dimensional embedding for each market node and product node. GNN is an effective framework that naturally operates on graph-structured data and learns relational information through message-passing between the nodes on the graph (Wu et al., 2021; Cappart et al., 2023). Therefore, we leverage GNNs in the input embedding module to embed the bipartite graph and capture the relations between markets and products.
Considering the feature heterogeneity of the bipartite graph representation, we first normalize the raw node and edge features by the product demands and linearly project them into the initial embeddings with a uniform dimension. Then, based on the structure of the input bipartite graph, we design a two-phase message-passing procedure that updates the initial node embeddings by aggregating information across the market nodes and product nodes: the first phase is performed to update the embeddings of product nodes, followed by the second phase that updates the embeddings of market nodes. Specifically, let $g_{k}^{\text{init}}$, $h_{m}^{\text{init}}$ and $e_{km}^{\text{init}}$ denote the initial embeddings of product node $k$, market node $m$, and edge $(k,m)$, respectively. The two-phase update is processed as follows:
@@
where $\operatorname{MLP}_{\mathcal{K}}$ and $\operatorname{MLP}_{\mathcal{M}}$ are multi-layer perceptrons (MLPs) for updating the product nodes and market nodes, respectively, and $\mathcal{N}\left(k\right)$ and $\mathcal{N}\left(m\right)$ denote the 1-hop neighbor node set of product node $k$ and market node $m$, respectively. Both phases can be viewed as a modification of the graph isomorphism network (GIN) (Xu et al., 2019) that further takes into account the bipartite structure and incorporates edge embeddings in the update. The updated product node embeddings $g_{k}^{0}$ for $k=1,2,\dots,|K|$ and market node embeddings $h_{m}^{0}$ for $m=0,1,\dots,|M|$ are then used in the market encoder and the decoder.
We remark that our design behind the two-phase embedding procedure is intuitive: in the first phase, we make each product node collect its supply information, including in which markets it is supplied, the coordinates of these markets, as well as the supply quantities and prices; and in the second phase, we make each market node collect higher-level information about its supplied products, including not only their total demands but also the information about how these products are supplied at other markets, which is extracted based on the product node embeddings from the first phase. Consequently, the product node embedding $g_{k}^{0}$ aggregates its global supply information, and the market node embedding $h_{m}^{0}$ collects more complex information such as the substitutive and complementary relations of product supply with other markets, which is important for the market selection decisions in route construction.
@@ 5.1.2 Market Encoder
It is empirically observed from prior works for routing problems that deeper information extraction for the city nodes, i.e., market nodes in our task, can potentially improve the capability of the policy network (Kwon et al., 2020; Joshi et al., 2022). Thus, we process the market node embeddings through a market encoder for further information extraction.
@@
where $h_{m}^{\ell+1}$ is the embedding of market node $m$ after layer $\ell$ for $\ell=0,\ldots,N-1$. The MHA layer adopts a scaled dot-product attention mechanism, where the attention function can be viewed as a mapping from a query vector and a set of key-value pairs to an output vector. Specifically, for each attention head, a query vector $q_{m}^{\ell}$, a key vector $k_{m}^{\ell}$, and a value vector $v_{m}^{\ell}$ are calculated for each market node $m$ through a linear projection of its node embedding:
@@
where $W_{q}^{\ell}$, $W_{k}^{\ell}$, and $W_{v}^{\ell}$ are trainable parameters. The attention output for market node $m$ is computed as a weighted sum of the values $v_{0}^{\ell},\ldots,v_{\left|M\right|}^{\ell}$, where the weight assigned to each value is the scaled dot-production of the query $q_{m}^{\ell}$ with the corresponding keys $k_{0}^{\ell},\ldots,k_{\left|M\right|}^{\ell}$:
@@
where $\text{dim}_{k}$ is the dimension of key vectors. Multi-head attention can be interpreted as multiple attention functions in parallel, which allows the node to collect diverse messages from other nodes through different attention heads. Then, we obtain the final MHA value of market node $m$ through a linear projection $W_{o}^{\ell}$ of the concatenated attention heads:
@@
where $\text{head}_{m,i}^{\ell}$ is an abbreviation for $\operatorname{Attention}^{\ell}_{i}(h_{m}^{\ell})$ and [,][$\cdot$,$\cdot$][ ⋅ , ⋅ ] is the concatenation operator.
After the $N$ attention layers, we compute the mean of final market node embeddings as a global embedding $h_{\mathcal{M}}$, which can be viewed as an aggregation of the final extracted information for the bipartite graph:
@@
The global embedding $h_{\mathcal{M}}$ is then used as part of the decoding context for the decoder.

### v6

@@ 3.2.3 Transition
where $q_{v^{t}k}$ is the available quantity of product $k$ supplied at market $v^{t}$. Otherwise, if the next node $v^{t}$ is the depot, a complete route is formed and the MDP terminates and then the reward calculation procedure for the terminal state is executed. Note that the TPP instance $U$, i.e., the static part of the states, remains fixed throughout the state transitions in an MDP.
@@ 3.2.4 Reward
After the sequential route construction process as described above, the second stage—purchase planning—is executed, which is integrated into the reward calculation procedure for the terminal state. Given the constructed route, a linear transportation problem is solved to derive the optimal purchasing plan for the visited markets. In this transportation problem, each source point corresponds to a market on the route, and each destination point corresponds to a product to be purchased. The unit transportation cost from source point $i$ to destination point $k$ is exactly the cost to purchase a unit of product $k$ from market $i$ in the TPP. This transportation problem can be efficiently solved using an off-the-shelf LP solver or a polynomial-time dynamic programming algorithm (Munkres, 1957). So far, a complete solution, including a route and the associated purchasing plan, has been obtained.
The objective of TPPs is to minimize the sum of traveling and purchasing costs while meeting all problem constraints. Since the feasibility of the route and the purchasing plan have been guaranteed by the masking rules (Section 3.2.2), we define the reward for the terminal state as the negative objective value, i.e., the sum of traveling and purchasing costs, and assign a zero-reward for all intermediate steps. The reward is used to update the policy network using a DRL training algorithm, which will be described in detail in Section 5.2. This way, the policy network is trained to optimize the global objective, that is, to construct a high-quality route that minimizes not only the traveling cost but also the subsequent purchasing cost based on it, thus to minimize the total cost.
@@ 3.2.5 Policy
Given the state $s_{t}$ at each step $t$, the action $a_{t}$ is decided based on a stochastic policy $p_{\theta}(a_{t}\mid s_{t})$, which can be viewed as a distribution over the action set $A_{t}$. Since the transition from a state $s_{t}$ to the next state $s_{t+1}$ is deterministic given an action $a_{t}$ (Section 3.2.3), the joint probability of producing a route $\pi$ based on the stochastic policy $p_{\theta}(a_{t}\mid s_{t})$ can be factorized according to the chain rule as:
@@
where the route is constructed in $T$ steps, and the action $a_{t}$ is sampled based on the policy $p_{\theta}(a_{t}\mid s_{t})$ at each step. In our DRL-based approach, the policy is parameterized as a deep neural network $\theta$, where the input of the policy network is the state $s_{t}$ and the output is the policy, i.e., the action distribution $p_{\theta}(a_{t}\mid s_{t})$. The DRL training aims to learn a policy that can produce high-reward solutions with high probabilities and low-reward solutions with low probabilities.
@@ 4 Bipartite Graph Representation for TPPs
Similar to existing DRL-based methods for routing problems (Kool et al., 2019; Joshi et al., 2019), a natural and straightforward idea is to represent a TPP instance as a complete or k-NN graph, where each market is represented as a node, connected to each other, and the product supply information at each market is encoded as node features. However, despite containing necessary information for the TPP instance, such representations impose many drawbacks when used in DRL. First, the feature dimensions vary with the number of products in the TPP instance being solved. Specifically, two features are needed for each product—the available quantity and the price. Consequently, a 50-product instance requires a 100-dimensional feature vector for each node to record the supply information for these 50 products, while an instance with 100 products would require 200 dimensions. This will lead to a computational mismatch in the forward-propagation of neural networks. In other words, a policy network can only be applied on fixed-sized TPP instances, unable to adapt to instances with different sizes even if only one additional product is introduced, significantly limiting its practical usability. Second, the relations between markets and products, such as the substitutions and complementarities in product supply between different markets, are very important for constructing a low-cost route. However, such dependencies are largely ignored in the complete or k-NN graph representations. To address these limitations, we design a novel bipartite graph representation, which is built upon the MILP formulation of TPPs.
@@
Bipartite graphs have been demonstrated effective in representing general MILP problems (Nair et al., 2020; Morabit et al., 2021), where the variables and constraints are represented as two sets of nodes, and an edge connects a variable node and a constraint node if the corresponding column contributes to this constraint. Readers are referred to Chen et al. (2023) for a more detailed introduction. However, directly applying this representation on the MILP formulation (1)-(9) for TPPs can introduce unnecessary redundancy. Specifically, the parameters to characterize a specific TPP instance appear only in the objective function (1) and constraints (2)-(3), whereas constraints (4)-(6), which enforce the formation of a simple cycle, are already guaranteed by the masking rules (Section 3.2.2) and thus need not be explicitly represented for the policy network. Moreover, the constraints and variables associated with the same element (e.g., a market or a product) can be merged to reduce the graph size and thus improve the computational efficiency of the policy network.
@@
Therefore, we design a bipartite graph representation for TPPs based on the structure of the objective function (1) and constraints (2)-(3), with appropriate merging of the constraints and variables. Our bipartite graph representation is shown as Figure 2. We represent the constraints associated only with products (i.e., constraints (2)) and the variables associated only with markets / depot (i.e., variables $v_{i}$ and $x_{ij}$) as two sets of nodes, and the constraints and variables associated with both products and markets (i.e., constraints (3) and variables $z_{ik}$) as edges connecting the associated product nodes and market / depot nodes. Specifically, each variable $v_{i}$ is represented as a market / depot node, with the objective coefficients $c_{ij}$ of variables $x_{ij}$ represented by coordinate features of the market / depot nodes. Each constraint in (2) is represented as a product node, with the right-hand side values $d_{k}$ attached as node features. An edge connects a market node $i$ and a product node $k$ if the associated variable $z_{ik}$ is not always restricted to 00 in constraints (3), that is, product $k$ is available at market $i$. The edge feature is a 2-dimensional vector, which consists of the corresponding objective coefficient $p_{ik}$ of variable $z_{ik}$ and the constraint coefficient $q_{ik}$ in constraints (3). For simplicity in the remainder of this paper, we use the term “market nodes” to refer to “market / depot nodes” if it does not cause confusion. We remark that our bipartite graph representation can be viewed from a more intuitive perspective: it captures the two fundamental elements of TPPs—markets and products—as two sets of nodes, while their interdependencies, such as the available quantities and prices, are characterized using the edges between them.
@@
We emphasize that our bipartite graph representation effectively addresses the limitations of the complete or k-NN graph representations discussed above. First, the dimensions of node and edge features are invariant to the number of markets and products, enabling the design of a size-agnostic policy network, such that it can be flexibly adapted to TPP instances of varying sizes. Second, the bipartite graph explicitly characterizes the relations between markets and products through edges connecting the market nodes and product nodes. The message-passing along edges can effectively capture the relational information between markets and products, which is important to the decision tasks.
@@ 5 Policy Network
In this section, we introduce the architecture of our policy network, which is used for route construction in the first stage. In addition, at the end of this section, we will describe the basic training algorithm for updating the policy network during training.
@@ 5.1 Policy Network Architecture
As described above, at each decision step $t$, the policy network, denoted as $\theta$, takes as input the state $s_{t}$ (consisting of the TPP instance $U$ being solved, which is represented as a bipartite graph, the partial route $\pi^{t}$, and the remaining demand $d^{t}$), and outputs a distribution $p_{\theta}(a_{t}|s_{t})$ over the actions $a_{t}\in A_{t}$, determining the next node to visit. The architecture of our policy network is illustrated in Figure 3. The policy network is composed of 1) an input embedding module, 2) a market encoder, and 3) a decoder.
The encoding process is designed to fully exploit the structure of bipartite graph. Specifically, the input embedding module takes the bipartite graph representation of instance $U$ as input, performing message-passing along edges to produce high-dimensional embeddings for the market and product nodes. These market node embeddings are then further processed through the market encoder to extract relevant information between each other. Following the encoding process, a decoder is executed iteratively to construct the route in a sequential manner. Specifically, at each decision step $t$, the decoder receives a decoding context that contains the embeddings of the bipartite graph and the dynamic part of current state $s_{t}$ (i.e., $\pi^{t}$ and $d^{t}$) as input. Based on this, the decoder outputs a distribution $p_{\theta}(a_{t}\mid s_{t})$, i.e., the policy, from which an action $a_{t}$ is selected, determining the next node to visit. The dynamic part $\pi^{t}$ and $d^{t}$ of the state are updated accordingly, starting the next decoding step. Since the bipartite graph representation of $U$ is fixed throughout the state transitions in a complete MDP, we execute the input embedding module and market encoder only once at the initial state, and then the decoder is executed iteratively with each decision step of the MDP. In the following of this section, we provide a detailed introduction to the three components of our policy network.
Figure 3: The architecture of our policy network. The policy network is composed of 1) an input embedding module, 2) a market encoder, and 3) a decoder. The bipartite graph is embedded through the input embedding module and the market encoder. The decoder is executed iteratively: at step $t$, the decoder receives a decoding context, and outputs an action distribution $p_{\theta}(a_{t}\mid s_{t})$, from which an action $a_{t}$ is selected, determining the next node to visit.
@@ 5.1.1 Input Embedding Module
Taking the bipartite graph representation of instance $U$ as input, the input embedding module exploits the raw features and topology of the bipartite graph and produces a high-dimensional embedding for each market node and product node. GNN is an effective framework that naturally operates on graph-structured data for learning relational information through message-passing between the nodes on the graph (Wu et al., 2021). Therefore, we incorporate GNNs into our input embedding module to embed the bipartite graph and capture the relations between markets and products.
Considering the feature heterogeneity of the bipartite graph representation, we first normalize the node and edge features by the product demand values, and linearly project them into a uniform-dimensional space as initial embeddings. To exploit the structure of the bipartite graph, we introduce a two-phase message-passing procedure that updates the initial node embeddings by aggregating information across the market nodes and product nodes: the first phase is performed to update the embeddings of product nodes, followed by the second phase that updates the embeddings of market nodes. Specifically, let $g_{k}^{\text{init}}$, $h_{m}^{\text{init}}$ and $e_{km}^{\text{init}}$ denote the initial embeddings of product node $k$, market node $m$, and edge $(k,m)$, respectively. The two-phase update proceeds as follows:
@@
where $\operatorname{MLP}_{\mathcal{K}}$ and $\operatorname{MLP}_{\mathcal{M}}$ are multi-layer perceptrons (MLPs) for updating the product nodes and market nodes, respectively, and $\mathcal{N}\left(k\right)$ and $\mathcal{N}\left(m\right)$ denote the 1-hop neighbor node set of product node $k$ and market node $m$, respectively. Both phases can be seen as an extension of graph isomorphism network (GIN) (Xu et al., 2019), adapted to further take the bipartite structure into account and incorporate edge embeddings in the update. The updated product node embeddings $g_{k}^{0}$ for $k=1,2,\dots,|K|$ and market node embeddings $h_{m}^{0}$ for $m=0,1,\dots,|M|$ are used in the downstream market encoder and the decoder.
We remark that our design behind the two-phase embedding procedure is intuitive. In the first phase, each product node gathers its supply information, including in which markets it is supplied, the coordinates of these markets, as well as the supply quantities and prices. This allows each product node to form an “understanding” of where and how it is supplied. In the second phase, we make each market node collect higher-level information about its supplied products, including not only their total demands but also the information about how these products are supplied at other markets. This information is derived from the product node embeddings updated in the first phase. Consequently, the product node embedding $g_{k}^{0}$ aggregates its global supply information, and the market node embedding $h_{m}^{0}$ incorporates higher-level information such as the substitutive and complementary relations of product supply with other markets, which is important for making informed market selection decisions.
@@ 5.1.2 Market Encoder
It is empirically observed from prior works for routing problems that deeper information extraction for the city nodes—i.e., market nodes in our task—can potentially improve the capability of the policy network (Kwon et al., 2020; Joshi et al., 2022). Thus, we further process the market node embeddings through a market encoder for deeper information extraction.
@@
where $h_{m}^{\ell+1}$ is the embedding of market node $m$ after layer $\ell$ for $\ell=0,\ldots,N-1$. The MHA layer adopts a scaled dot-product attention mechanism, mapping a query vector and a set of key-value pairs to an output vector. For each attention head, a query vector $q_{m}^{\ell}$, a key vector $k_{m}^{\ell}$, and a value vector $v_{m}^{\ell}$ are calculated for each market node $m$ through a linear projection of its node embedding:
@@
where $W_{q}^{\ell}$, $W_{k}^{\ell}$, and $W_{v}^{\ell}$ are trainable parameters. The attention output for market node $m$ is calculated as a weighted sum of the values $v_{0}^{\ell},\ldots,v_{\left|M\right|}^{\ell}$, where the weight assigned to each value is the scaled dot-production of the query $q_{m}^{\ell}$ with the corresponding keys $k_{0}^{\ell},\ldots,k_{\left|M\right|}^{\ell}$:
@@
where $\text{dim}_{k}$ denotes the dimension of key vectors. Multi-head attention can be interpreted as multiple attention functions in parallel, which enables the node to collect diverse messages from other nodes through different attention heads. The final MHA output for each market node $m$ is calculated by linearly projecting the concatenation of the attention heads:
@@
where $\text{head}_{m,i}^{\ell}$ is an abbreviation for $\operatorname{Attention}^{\ell}_{i}(h_{m}^{\ell})$ and [,][$\cdot$,$\cdot$][ ⋅ , ⋅ ] denotes the concatenation operator.
After passing through $N$ attention layers, we calculate the mean of final market node embeddings as a global embedding $h_{\mathcal{M}}$, which is an aggregation of the final extracted information for the bipartite graph:
@@
The global embedding $h_{\mathcal{M}}$ is then used as part of the decoding context, allowing the decoder to make well-informed decisions based on the global information of the instance.